OpenStudy (anonymous):

olve for x: 3(x + 1) = -2(x - 1) + 6.

OpenStudy (redheadangel):

First distribute

OpenStudy (redheadangel):

then move the variable on the right to the left by subtracting

OpenStudy (redheadangel):

then subtract 4 to the right side

OpenStudy (redheadangel):

then divide, then : BOOM YOUR ANSWER!

OpenStudy (anonymous):

3x + 3= -2x+ 6 3= -5x +3 6= -5x -6/5=x <=== this is the answer